OpenStudy (anonymous):

g(x)=X^2-2x Find g(x-1) Please explain to me how to plug in everything.

OpenStudy (anonymous):

I want to know how you get the answer step by step.

zepdrix (zepdrix):

Hey :) So we have g being a function of x. That means that you put some value in place of all the x's, and you get g as your output.\[\large\rm g(\color{orangered}{x})=(\color{orangered}{x})^2-2(\color{orangered}{x})\]

zepdrix (zepdrix):

Here is a quick example. For x=3, our function would look like this:\[\large\rm g(\color{orangered}{3})=(\color{orangered}{3})^2-2(\color{orangered}{3})\]Simplifying, we would determine that\[\large\rm g(3)=3\] Because 9-6=3 k? :)

zepdrix (zepdrix):

So instead of plugging in a numerical value for x, we're going to plug in x-1, it might seem a little strange at first.

zepdrix (zepdrix):

\[\large\rm g(\color{orangered}{x-1})=(\color{orangered}{x-1})^2-2(\color{orangered}{x-1})\]

zepdrix (zepdrix):

And then we need to expand out the right side, and combine like-terms and all that.

zepdrix (zepdrix):

Making ... a little sense up to this point? Do you understand how to expand out this square?\[\large\rm (x-1)^2=?\]

OpenStudy (anonymous):

yes it would be (x-1) (x-1)

OpenStudy (anonymous):

which would be x^2-2x+1

zepdrix (zepdrix):

\[\large\rm g(x-1)=x^2-2x+1-2(x-1)\]Good! Expand out the other term, and combine stuff.

OpenStudy (anonymous):

so is the answer be x^2-4x+3?

OpenStudy (anonymous):

if my answer is correct then I am good to go!

zepdrix (zepdrix):

Yes! Good job!